OpenStudy (anonymous):

Find the zeros of the function: f(x) = 3x3 - 12x2 - 15x

OpenStudy (johnweldon1993):

what can be factored out of each of the terms?

OpenStudy (johnweldon1993):

also notice that there IS at least 1 "x" in each term.....so you can factor out an x as well

OpenStudy (anonymous):

First use the factor theorem, by which we know that the factors of -15 divided by the factors of 3 are potential 0s. List all these out and try some. As soon as you find one that results in f(x) = 0, you know that the value is a 0. For example, if 1/2 is a potential zero of an arbitrary third degree function and by inserting x = 1/2 f(x) = 0, then 2x - 1 is a zero. Dividing the expression by 2x - 1 using long division will then result in a quadratic which you can normally factor and get all the zeroes. Try doing this to find the the zeros of this third degree polynomial. @nadc2005

OpenStudy (anonymous):

\[3x ^{3}-12x ^{2}-15x = 0\] Factor the left side, \[3(x-5)x(x+1)=0\] \[(x-5)x(x+1)=0\] We take the 3 to the other side and divide it by 0, which gives us 0. Then equate each part to 0 individually, \[x=0\] \[x-5 = 0; x = 5\] \[x+1 = 0; x=-1\] So the zeroes are 0, 5, and -1.